# Symmetry Breaking and Spatiotemporal Pattern Formation in Photonic Time Crystals

### v1

@@ Abstract
In this work, we explore the dynamics of time varying photonic media with an optical Kerr nonlinearity and an associated phase transition. The interplay between a periodically modulated permittivity and the nonlinearity induces a continuous transition of electromagnetic waves to a state with broken spatial and time translation symmetries. This transition gives rise to a lattice-like wave pattern, in many ways similar to a spatial crystallization in solids. Symmetry breaking triggers the emergence of soft, Goldstone-like modes, which propagate as deformations of the lattice structure, as well as massive Higgs-like modes – spatially uniform oscillations of the field amplitude. We extend the analysis of the non-equlibrium symmetry breaking to 2+1 dimensional time varying media and discuss pattern formation as well as the connection to discrete time crystals.
@@ I Introduction
In this paper, we study nonlinear photonic time crystals (nonlinear PTCs) – dielectric media with a periodically modulated permittivity [22, 23]. It is crucial to distinguish PTCs from “time crystals” in the sense of Wilczek [24, 25, 26, 27, 28]. While it is important to maintain this distinction, we demonstrate that, in the presence of a Kerr term, a PTC evolves to a subharmonic steady state oscillating at half the driving frequency, and thus breaks the discrete time translation symmetry of the drive on long timescales. The steady state is very much akin to Faraday waves [29, 30, 31], which are in many ways similar to classical discrete time crystals [32, 33, 28].
@@ II Results
Figure 1: a) Real part of the dispersion relation $\omega\left(q\right)$ of the PTC with momentum gaps at $q^{*}=\Omega/c$, where $2\Omega$ is the modulation frequency of the dielectric constant. b) Effective potential $V\left(\bar{b}\right)$ for the displacement field amplitude $\bar{b}$ with minima corresponding to steady states with broken translational and discrete time translational symmetries (see Eq. (9)). Here, $\epsilon$ is a parameter that encodes the relative strength of the modulation of the dielectric constant with respect to damping (see Eq. (5)). c) A numerical simulation of Eq. (1) showing the self organization of electric fields into the symmetry breaking steady state, corresponding to a minimum of the potential $V\left(\bar{b}\right)$. To produce the plot, we choose a modulation strength close to the threshold yielding $\epsilon\approx 0.02$. The $t$- and $x$-axes are drawn in units of $T=2\pi/\Omega$ and $\lambda=2\pi/q^{*}$.
@@
where $D\left(x,t\right)$ is the displacement field, $\gamma$ is the decay rate of the electromagnetic field, $c$ is the speed of light and $h$ is a dimensionless parameter controlling the modulation strength of the dielectric constant. In what follows we assume $h>0$. For brevity, we omit vector signs. For a small $h\ll 1$, an approximate solution to Eq. (1) can be found with the slowly varying envelope approximation. Anticipating that the time dependent term will lead to a parametric response at half the modulation frequency $\Omega$ [38], we choose the ansatz
@@
Let us now turn to the weakly damped regime where the modulation strength $h$ is close to the instability threshold $h_{c}$. This regime is useful to study the transition dynamics from the trivial to the symmetry breaking state. In particular, it can be shown that the transition into the symmetry breaking steady state is a continuous phase transition. To this purpose, we rewrite the amplitude equations for $a\left(t\right)$, $b\left(t\right)$ in therm so the new variables $\bar{a}\left(t\right)=a\left(t\right)+b\left(t\right)$, $\bar{b}\left(t\right)=a\left(t\right)-b\left(t\right)$. A perturbative expansion of Eqs. (4) in the small parameter $\epsilon$ of Eq. (5) (see supplement) then predicts that the transitional dynamics can be described by a gradient descent equation for $\bar{b}$$\left(t\right)$:
@@
We plot the effective potential for different $\epsilon$ in Fig. 1b and conclude that the transition into the symmetry breaking state is continuous, i.e. the system goes through a soft bifurcation when $h$ reaches the critical driving strength $h_{c}$. In terms of the original $a$ and $b$, the potential minima (and approximate fixed points of Eqs. (4)) are located at
$a_{0}=-b_{0}=\pm\sqrt{\frac{\tilde{h}_{c}}{\tilde{\beta}}}\left(2\epsilon\right)^{1/4}.$ (10)
@@
Figure 3: In two dimensions, nonlinear PTCs evolve to a stripe state that breaks the translational and rotational symmetries of the system. The figure shows the simulation of two-dimensional non-linear PTC starting from random initial conditions on square domain with commensurate periodic boundary conditions. First, the system develops inhomogeneities whose length scale is given by the critical $q^{*}$. Finally, a two dimensional stripe pattern, as predicted in Eq. (15), appears (see supplement for detailed figures).
@@ III Discussion
Summing up, we investigated the long term behavior of a photonic time crystal with a generic Kerr nonlinearity. We have shown that the initially exponential growth of unstable gap modes in these systems heralds the transition to a new symmetry breaking steady state. In this state, the field configuration is a standing wave lattice, where the lattice spacing is determined by the resonant wavenumber $q^{*}$, fulfilling the resonance condition $\Omega=cq^{*}$. The breaking of continuous spatial translation symmetry is accompanied by the appearance of soft Goldstone-like modes, which consist of propagating, wavelike lattice distortions, similar to phonons in a crystal lattice. In addition, gapped modes that correspond to amplitude oscillations of the electric field emerge. The dispersions of these emergent modes depend on the driving strength, which opens the possibility to use nonlinear PTCs in the symmetry broken state as tunable metamaterials. We also investigated the behavior of nonlinear PTCs in two spatial dimensions, and showed that the electric fields self-organize to stripe patterns that break the translational, and rotational symmetries of the system. This behavior could help to create controllable, spatially structured high intensity electric fields.
Furthermore, the predicted steady state exhibits subharmonic response, and breaks the discrete time translation symmetry of the periodic modulation. This is a common feature of many types of parametrically driven nonlinear waves [43, 31], as it has been pointed out in recent literature on time crystals [32, 33, 28]. Thus the nonlinear PTC exhibits discrete time crystalline behavior in the sense of a phase of matter [28].
@@ Acknowledgements.
Supplementary Material
Supplementary Sec..1 Expansion close to the instability threshold. Continuity of the phase transition.
@@
$\dot{\bar{b}}$ $=\Omega\tilde{h}_{c}\epsilon\bar{b}-\frac{\tilde{\beta}\Omega}{2}\bar{a}\left(\bar{a}^{2}+\bar{b}^{2}\right).$ (S 1)
@@
$\bar{a}_{2}$ $=-\frac{\tilde{\beta}}{4\tilde{h_{c}}}\bar{b}_{1}^{3}.$ (S 2)
@@
For a small $\epsilon$, we can assume that the relaxation of $\bar{a}_{2}$ towards the value given in Eq. (S 2) is infinitely fast, and write
$\frac{\partial\bar{b}_{1}}{\partial t_{1}}=\tilde{h}_{c}\Omega\bar{b}_{1}-\frac{\tilde{\beta}^{2}\Omega}{8\tilde{h_{c}}}\bar{b}_{1}^{5}.$ (S 3)
@@
Supplementary Sec..2 Goldstone-like Modes
@@
$D_{\phi}\left(x,t\right)=\pm\sqrt{\frac{8h}{9\beta}}\sin\left(\Omega t\right)\cos\left(q^{*}x+\phi\left(x,t\right)\right).$ (S 4)
@@
$\quad+\frac{A}{2}\sin\left(\Omega t+qx+\phi\left(t,x\right)\right),$ (S 5)
where $A=\sqrt{8h/9\beta}$. In what follows, we will use the abbreviations $\sin\left(\Omega t-qx-\phi\left(t,x\right)\right)=\sin\left(-\right)$ and $\sin\left(\Omega t+qx+\phi\left(t,x\right)\right)=\sin\left(+\right)$ and similar for $\cos$-terms. Inserting Eq. (S 5) into Eq. (1) and comparing the coefficients in front of the $\cos\left(+\right)$ terms we find an equation for $\phi\left(t,x\right)$:
$\frac{1}{c^{2}}\frac{\partial^{2}\phi}{\partial t^{2}}=\left(1-\frac{1}{3}h\right)\frac{\partial^{2}\phi}{\partial x^{2}}.$ (S 6)
The same equation follows from the $\cos\left(-\right)$ terms. The $\sin\left(\pm\right)$ terms, on the other hand, give corrections of order $Q/q^{*}$ to the solution (S 4). These can be neglected for a smooth, small-$Q$ perturbation of the kind that we are considering. Compared to the velocity of electromagnetic waves in the undriven medium, the velocity of the Goldstone mode $c_{G}$ is reduced by a factor that depends on the amplitude of the time varying term in Eq. (1):
@@
Supplementary Sec..3 Pattern formation in two dimensions
Supplementary Figure S 1: Pattern formation in a two-dimensional PTC.

### v2

@@ Abstract
In this work, we explore the dynamics of time varying photonic media with an optical Kerr nonlinearity and an associated phase transition. The interplay between a periodically modulated permittivity and the nonlinearity induces a continuous transition of electromagnetic waves to a state with broken spatial and time translation symmetries. This transition gives rise to a lattice-like wave pattern, in many ways similar to a spatial crystallization in solids. Symmetry breaking triggers the emergence of soft, Goldstone-like modes, which propagate as deformations of the lattice structure, as well as massive Higgs-like modes – spatially uniform oscillations of the field amplitude. We extend the analysis of the non-equlibrium symmetry breaking to 2+1 dimensional time varying media and discuss pattern formation as well as the connection to discrete dissipative time crystals.
@@ I Introduction
In this paper, we study nonlinear photonic time crystals (nonlinear PTCs) – dielectric media with a periodically modulated permittivity [22, 23]. It is crucial to distinguish PTCs from “time crystals” and “discrete time crystals” in the sense of Wilczek [24, 25, 26], which have been intensively studied in theory [27, 28] and experiment [29, 30, 31, 32, 33, 34, 35]. While it is important to maintain this distinction, we demonstrate that, in the presence of a Kerr term, a PTC evolves to a subharmonic steady state oscillating at half the driving frequency, and thus breaks the discrete time translation symmetry of the drive on long timescales. The steady state is very much akin to Faraday waves [36, 37, 38], which are in many ways similar to classical discrete time crystals [39, 40, 26]. In fact, an observation of Faraday waves in a cold atom Bose-Einstein-Condensate (BEC) [41] has recently been reinterpreted as experimental evidence for a discrete quantum space-time crystal [42, 43, 44, 45, 46]. An axial mode of a BEC trapped in a cigar-like potential was parametrically excited by a long-lived, radial oscillatory mode. This led to the formation of a periodic spatial structure in the axial direction.
@@
PTCs have already been realized as metamaterials in the GHz range [5]. Plasmonic realizations that are within experimental reach and could extend the frequency range to THz have been proposed [51, 52, 53, 54]. A promising route to realizing PTCs at optical frequencies are epsilon near zero materials [7] where the time refraction of optical signals has recently been demonstrated [55] – a promising milestone on the way to creating PTCs at optical frequencies.
@@ II Results
Figure 1: a) Real part of the dispersion relation $\omega\left(q\right)$ of the PTC with momentum gaps at $q^{*}=\Omega/c$, where $2\Omega$ is the modulation frequency of the dielectric constant. b) Effective potential $V\left(\bar{b}\right)$ for the displacement field amplitude $\bar{b}$ with minima corresponding to steady states with broken translational and discrete time translational symmetries (see Eq. (10)). Here, $\epsilon$ is a parameter that encodes the relative strength of the modulation of the dielectric constant with respect to damping (see Eq. (6)). c) A numerical solution of Eq. (1). Starting from a random field distribution at $t=0$, the plot shows the spatio-temporal evolution of the field $D\left(x,t\right)$ to the symmetry breaking state (checkerboard-like pattern) given in Eq. (2) and corresponding to a minimum of the potential $V\left(\bar{b}\right)$. To produce the plot, we choose a modulation strength close to the threshold yielding $\epsilon\approx 0.02$. The $t$- and $x$-axes are drawn in units of $T=2\pi/\Omega$ and $\lambda=2\pi/q^{*}$.
@@
where $D\left(x,t\right)$ is the displacement field, $\gamma$ is the decay rate of the electromagnetic field, $c$ is the speed of light and $h$ is a dimensionless parameter controlling the modulation strength of the dielectric constant. In what follows we assume $h>0$. For brevity, we omit vector signs. For a small $h\ll 1$, an approximate solution to Eq. (1) can be found with the slowly varying envelope approximation. From the condition of parametric resonance [56], we know that an oscillating system is parametrically excited when the modulation frequency is equal to twice the resonance frequency divided by an integer $n$. The $n=1$ resonance is the most pronounced. Correspondingly, we anticipate that the time dependent term will lead to a parametric response at half the modulation frequency $\Omega$ [56], and we choose the ansatz
@@
Let us now turn to the weakly damped regime where the modulation strength $h$ is close to the instability threshold $h_{c}$. This regime is useful to study the transition dynamics from the trivial to the symmetry breaking state. In particular, it can be shown that the transition into the symmetry breaking steady state is a continuous phase transition. To this purpose, we rewrite the amplitude equations for $a\left(t\right)$, $b\left(t\right)$ in terms of the new variables $\bar{a}\left(t\right)=a\left(t\right)+b\left(t\right)$, $\bar{b}\left(t\right)=a\left(t\right)-b\left(t\right)$. A perturbative expansion of Eqs. (4) in the small parameter $\epsilon$ of Eq. (6) (see supplement) then predicts that the transitional dynamics can be described by a gradient descent equation for $\bar{b}$$\left(t\right)$:
@@
We plot the effective potential for different $\epsilon$ in Fig. 1b and conclude that the transition into the symmetry breaking state is continuous, i.e. the system goes through a soft bifurcation when $h$ reaches the critical driving strength $h_{c}$, which is determined by the ratio of $\gamma$ and $\Omega$ (see Eq. (5)). In terms of the original $a$ and $b$, the potential minima (and approximate fixed points of Eqs. (4)) are located at
$a_{0}=-b_{0}=\pm\sqrt{\frac{\tilde{h}_{c}}{\tilde{\beta}}}\left(\frac{\epsilon}{2}\right)^{1/4}.$ (11)
@@
Figure 3: In two dimensions, nonlinear PTCs evolve to a stripe state that breaks the translational and rotational symmetries of the system. The figure shows the simulation of two-dimensional non-linear PTC starting from random initial conditions on square domain with commensurate periodic boundary conditions. First, the system develops inhomogeneities whose length scale is given by the critical $q^{*}$. Finally, after a series of complex transient patterns, a two dimensional stripe pattern, as predicted in Eq. (16), appears.
@@ III Discussion
We investigated the long term behavior of a photonic time crystal with a generic Kerr nonlinearity. We have shown that the initially exponential growth of unstable gap modes in these systems heralds the transition to a new symmetry breaking steady state. In this state, the field configuration is a standing wave lattice, where the lattice spacing is determined by the resonant wavenumber $q^{*}$, fulfilling the resonance condition $\Omega=cq^{*}$. The predicted PTC steady state exhibits subharmonic response, and breaks the discrete time translation symmetry of the periodic modulation. This is a common feature of many types of parametrically driven nonlinear waves [61, 38], as it has been pointed out in recent literature on time crystals [39, 40, 26]. Thus the nonlinear PTC exhibits discrete dissipative time crystalline behavior in the sense of a phase of matter [26].
The breaking of continuous spatial translation symmetry is accompanied by the appearance of soft Goldstone-like modes, which consist of propagating, wavelike lattice distortions, similar to phonons in a crystal lattice. In addition, gapped modes that correspond to amplitude oscillations of the electric field emerge. The dispersions of these emergent modes depend on the driving strength, which opens the possibility to use nonlinear PTCs in the symmetry broken state as tunable metamaterials. We also investigated the behavior of nonlinear PTCs in two spatial dimensions, and showed that the electric fields self-organize to stripe patterns that break the translational, and rotational symmetries of the system. This behavior could help to create controllable, spatially structured high intensity electric fields. Goldstone-like and gapped modes as the ones predicted here can be expected to exist in other parametrically driven, spatially extended classical and quantum systems. Apart from Faraday waves in classical liquids, where collective behavior has been observed [62, 63], important candidates are discrete space-time crystals in trapped BECs [42, 43, 44, 45, 46].
@@ Acknowledgements.
Appendix
Appendix.1 Expansion close to the instability threshold. Continuity of the phase transition.
@@
$\dot{\bar{b}}$ $=\Omega\tilde{h}_{c}\epsilon\bar{b}-\frac{\tilde{\beta}\Omega}{2}\bar{a}\left(\bar{a}^{2}+\bar{b}^{2}\right).$ (A 1)
@@
$\bar{a}_{2}$ $=-\frac{\tilde{\beta}}{4\tilde{h_{c}}}\bar{b}_{1}^{3}.$ (A 2)
@@
For a small $\epsilon$, we can assume that the relaxation of $\bar{a}_{2}$ towards the value given in Eq. (A 2) is infinitely fast, and write
$\frac{\partial\bar{b}_{1}}{\partial t_{1}}=\tilde{h}_{c}\Omega\bar{b}_{1}-\frac{\tilde{\beta}^{2}\Omega}{8\tilde{h_{c}}}\bar{b}_{1}^{5}.$ (A 3)
@@
Appendix.2 Goldstone-like Modes
@@
$D_{\phi}\left(x,t\right)=\pm\sqrt{\frac{8h}{9\beta}}\sin\left(\Omega t\right)\cos\left(q^{*}x+\phi\left(x,t\right)\right).$ (B 4)
@@
$\quad+\frac{A}{2}\sin\left(\Omega t+qx+\phi\left(t,x\right)\right),$ (B 5)
where $A=\sqrt{8h/9\beta}$. In what follows, we will use the abbreviations $\sin\left(\Omega t-qx-\phi\left(t,x\right)\right)=\sin\left(-\right)$ and $\sin\left(\Omega t+qx+\phi\left(t,x\right)\right)=\sin\left(+\right)$ and similar for $\cos$-terms. Inserting Eq. (B 5) into Eq. (1) and comparing the coefficients in front of the $\cos\left(+\right)$ terms we find an equation for $\phi\left(t,x\right)$:
$\frac{1}{c^{2}}\frac{\partial^{2}\phi}{\partial t^{2}}=\left(1-\frac{1}{3}h\right)\frac{\partial^{2}\phi}{\partial x^{2}}.$ (B 6)
The same equation follows from the $\cos\left(-\right)$ terms. The $\sin\left(\pm\right)$ terms, on the other hand, give corrections of order $Q/q^{*}$ to the solution (B 4). These can be neglected for a smooth, small-$Q$ perturbation of the kind that we are considering. Compared to the velocity of electromagnetic waves in the undriven medium, the velocity of the Goldstone mode $c_{G}$ is reduced by a factor that depends on the amplitude of the time varying term in Eq. (1):